ann has 3 times as many beads as marie, ann gives marie 55 beads, now ann has half as many beads as marie. how many beads are all together?
Answer by ankor@dixie-net.com(15622)

ann has 3 times as many beads as marie, ann gives marie 55 beads,
now ann has half as many beads as marie.
how many beads are all together?
:
a = 3m
Gives 55, now
(a-55) for Ann
(m+55) for Marie
:
"now ann has half as many beads as marie."
(a-55) = .5(m+55)
a - 55 = .5m + 27.5
a = .5m + 27.5 + 55
a = .5m + 82.5
replace a with 3m
3m = .5m + 82.5
3m - .5 = 62.5
2.5m = 82.5
m = 82.5/2.5
m = 33 beads originally for Marie
and
3(33) = 99 beads originally for Ann
then
33 + 99 = 132 total beads